OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line shown in the graph. Please show all of your work. (-5,-2) (5,2)

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

delta(y)/delta(x)

OpenStudy (anonymous):

slope intercept form is y=mx+b, where m is the slope and b is the y intercept. to find m we take y2-y1/x2-x1 (change in y over change in x) => 2--2/5--5 => 4/10 => 2/5=m the y intercept is when x=0. At this point we have y=2/5 x + b ... im guessing you can see the where the line crosses y and x=0 in your graph :] just plug that number in for b

OpenStudy (anonymous):

lulu you have to work too . dont expect others to finish all work if you dont understand pose more questions